OpenStudy (anonymous):

FAN AND MEDAL VERY QUICK 5 MIN QUESTION!!!!!

OpenStudy (anonymous):

f(x) = x^2 + 4 g(x) = x + 4

OpenStudy (anonymous):

f(x) = g(x) Solve algebraically

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

...

OpenStudy (bibby):

\(x^2 + 4=x + 4\) solve

OpenStudy (anonymous):

Ok one sec

OpenStudy (anonymous):

Given f(x) = 2x + 3 and g(x) = –x2 + 5, find (g o f )(1).

OpenStudy (anonymous):

i got x^2 = x ???

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

exactly. now subtract x from both sides and factor

OpenStudy (anonymous):

I can factor but I don't know what you mean by subtract @bibby

OpenStudy (bibby):

subtract x from both sides

OpenStudy (anonymous):

x^2 - x = 0

OpenStudy (bibby):

if we're to solve for x, we need all the x's on one side of the equation

OpenStudy (bibby):

factor that

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

-1(-x^2 - x)

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@bibby